# April 2026

### 05 Apr 2026 v1

@@ 7 Conclusions and outlook
An important aspect concerns the analysis of the role of energy positivity. While the stress–energy tensor fails to define positive operators on the full Fock space due to the Reeh–Schlieder theorem, we have shown that suitable lower bounds can be established by means of quantum energy inequalities. This analysis makes it possible to control deviations from positivity and to define regularized families of positive operators $A^{u}_{f,\epsilon}$ that approximate the localization effects with arbitrary precision on states with a fixed particle number. The price to pay is, of course, that the operators $A^{u}_{f,\epsilon}$ are non-local; however, they arise as restrictions of non-local and non-positive operators ${\mathsf{A}}^{u}_{f,\epsilon}$ acting on the full Fock space.
@@ Appendix B Appendix: Proofs of technical propositions
The action of an operator $\partial^{|\alpha|}_{k^{\alpha}}$ on $\phi$ produces a linear combination of functions as in (115) with the following changes in the integrand: (a) factors $x^{\beta}$, (b) factors given by products of $k$-derivatives (of order $1$ at least) of the function $E(k)$. The terms of type (a) can be accommodated into a new definition of the Schwartz function $f^{\prime}$. We end up with a linear combination of functions
@@
Proof of Proposition 4.4. According to Propositions 3.5 and 3.9, the map $\mathscr{D}({\mathbb{M}})\ni g\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle=\langle\hat{\phi}[\overline{f}]\Psi|\hat{\phi}[g]\Psi^{\prime}\rangle-\langle\hat{\phi}[\overline{f}]\Omega|\hat{\phi}[g]\Omega\rangle\langle\Psi|\Psi^{\prime}\rangle$ is continuous for every given $f\in\mathscr{D}({\mathbb{M}})$. With a similar argument, $\langle\Psi|:\thinspace\hat{\phi}[f_{n}]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle\to 0$ if $f_{n}\to 0$ in $\mathscr{D}({\mathbb{M}})$ for every given $g\in\mathscr{D}({\mathbb{M}})$. Schwartz’ kernel theorem implies that there exists a unique distribution in $\mathscr{D}^{\prime}({\mathbb{M}}\times{\mathbb{M}})$ which extends the map $\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle$. We denote its kernel by $\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle$. At this juncture, using the definition (19), (44), and (107) one finds that the bounded smooth $x$-$y$ symmetric function (as the quadratic forms $(p,k)\mapsto\langle\Psi^{\prime}|a^{\#}_{p}a^{\#}_{k}\Psi\rangle$ are Schwartz functions on $\mathsf{V}_{m,+}^{2}$)
@@
where $p\equiv(E(p),\vec{p})$, $k\equiv(E(k),\vec{k})$ with $E(p)=p^{0}=\sqrt{\vec{p}^{2}+m^{2}}$ and $E(k)=k^{0}=\sqrt{\vec{k}^{2}+m^{2}}$. The integrated functions of $\vec{p}$ and $\vec{k}$ are of Schwartz type in ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\equiv\mathsf{V}_{m,+}^{2}$ by construction. Thus, in particular, one can pass the $x$-derivatives under the symbol of integration. Let us discuss the proof of (d). First of all we observe that the integrals on both sides are finite since the integrated functions are Schwartz, in view of the bound (55), which is valid in two Minkowskian coordinate systems comoving with the reference frames $u_{\Sigma}$ and $u_{\Sigma^{\prime}}$ respectively. Define $J^{\nu}:=v_{\mu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle$. We observe that $J$ is smooth and conserved, $\partial_{\mu}J^{\mu}(x)=0$, by (c). Now consider the case where $\Sigma$ and $\Sigma^{\prime}$ are parallel. In this case we can use a common system of Minkowskian coordinates, where $\Sigma$ is the plane $x^{0}=0$ and $\Sigma^{\prime}$ is the analogous plane at $x^{0}=T>0$ (or vice versa). If $C_{r}\subset{\mathbb{R}}^{4}$ is the cylinder with axis $\vec{x}=0$ and basis $B_{r}:=\{\vec{x}\in{\mathbb{R}}^{3}\>|\>|\vec{x}|<r\}$ in $\Sigma$, we can use the divergence theorem for the compact cylinder $G_{r}\subset C_{r}$ bounded by $B_{r}$ and the analogous basis $B^{\prime}_{r}$ in $\Sigma^{\prime}$. We move on to compute the flux of $J$ across $\partial G_{r}$. This flux must be 0 since $J$ is conserved, taking the divergence theorem into account. The flux of $J$ on the lateral surface of $G_{r}$ is bounded by $cTr^{2}$, so that it tends to 0 as $r\to+\infty$ due to (55). The outward fluxes on the bases respectively tend to $\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma}d\Sigma$ and $\int_{\Sigma^{\prime}}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma^{\prime}}d\mu_{\Sigma^{\prime}}$ in view of the dominated convergence theorem. This proves the thesis for the case under consideration. It remains to consider the case where $\Sigma$ and $\Sigma^{\prime}$ are not parallel and thus meet in a $2$-plane. Let us arrange a Minkowskian reference frame with the origin contained in this intersection and $\Sigma$ described by the plane $x^{0}=0$ in ${\mathbb{R}}^{4}$. With a further spatial rotation of the coordinates if necessary, we can describe the $3$-plane $\Sigma^{\prime}$ as the plane $x^{0}=\gamma x^{1}$ in ${\mathbb{R}}^{4}$ where $\gamma\in(0,1)$, since $\Sigma^{\prime}$ is spacelike. We assume that we deal with vectors $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ such that their (finite smooth) components $\Psi^{n},\Psi^{\prime n}$ have compact support in the corresponding spaces ${\mathbb{R}}^{3n}\equiv\mathsf{V}_{m,+}^{n}$. At this juncture, using the same proof of (b) in (19) of Theorem [CDM26] for the integrals (118), we have that the second bound in (b) is satisfied:
@@
Proof of Lemma 6.3. It is sufficient to apply Eq.(3) in [Few12] referring to the explicit computations presented in Sec.2.5 therein, specialized to the Minkowski spacetime and using as reference state the Poincaré invariant state we indicated by $\Omega$. (The only necessary hypothesis is that the two-point function of the considered normalized state vector $\Psi$ is such that ${\mathbb{M}}\times{\mathbb{M}}\ni(x,y)\mapsto\langle\Psi|\hat{\phi}(x)\hat{\phi}(y)\Psi\rangle-\langle\Omega|\hat{\phi}(x)\hat{\phi}(y)\Omega\rangle\in C^{\infty}({\mathbb{M}}\times{\mathbb{M}})$. For $\Psi\in{\mathfrak{S}}_{0}$ this fact is guaranteed by Proposition 4.4.) In this case the partial differential operator $Q$ is the operator $u^{\mu}u^{\prime\nu}D_{\mu\nu}$ (47) treated as in the proof of Proposition 4.10: $u^{\mu}u^{\prime\nu}D_{\mu\nu}=c^{2}D^{\prime}_{00}-s^{2}D^{\prime}_{11}$

### 07 Apr 2026 v2

@@ 7 Conclusions and outlook
An important aspect concerns the analysis of the role of energy positivity. While the stress–energy tensor fails to define positive operators on the full Fock space due to the Reeh–Schlieder theorem, we have shown that suitable lower bounds can be established by means of quantum energy inequalities. This analysis makes it possible to control deviations from positivity and to define regularized families of positive operators $A^{u}_{f,\epsilon}$ that approximate the localization effects with arbitrary precision on states with a fixed particle number. The price to pay is, of course, that the operators $A^{u}_{f,\epsilon}$ are non-local; however, they arise as restrictions (more precisely as compressions to the $n$-particle spaces) of non-local, bounded from below but non-positive operators ${\mathsf{A}}^{u}_{f,\epsilon}$ acting on the full Fock space.
@@ Appendix B Appendix: Proofs of technical propositions
The action of an operator $\partial^{|\alpha|}_{k^{\alpha}}$ on $\phi$ produces a linear combination of functions as in (115) with the following changes in the integrand: (a) factors $x^{\beta}$, (b) factors given by products of $k$-derivatives (of order 11 at least) of the function $E(k)$. The terms of type (a) can be accommodated into a new definition of the Schwartz function $f^{\prime}$. We end up with a linear combination of functions
@@
Proof of Proposition $4.$ According to Propositions 3.5 and 3.9, the map $\mathscr{D}({\mathbb{M}})\ni g\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle=\langle\hat{\phi}[\overline{f}]\Psi|\hat{\phi}[g]\Psi^{\prime}\rangle-\langle\hat{\phi}[\overline{f}]\Omega|\hat{\phi}[g]\Omega\rangle\langle\Psi|\Psi^{\prime}\rangle$ is continuous for every given $f\in\mathscr{D}({\mathbb{M}})$. With a similar argument, $\langle\Psi|:\thinspace\hat{\phi}[f_{n}]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle\to 0$ if $f_{n}\to 0$ in $\mathscr{D}({\mathbb{M}})$ for every given $g\in\mathscr{D}({\mathbb{M}})$. Schwartz’ kernel theorem implies that there exists a unique distribution in $\mathscr{D}^{\prime}({\mathbb{M}}\times{\mathbb{M}})$ which extends the map $\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle$. We denote its kernel by $\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle$. At this juncture, using the definition (19), (44), and (107) one finds that the bounded smooth $x$-$y$ symmetric function (as the quadratic forms $(p,k)\mapsto\langle\Psi^{\prime}|a^{\#}_{p}a^{\#}_{k}\Psi\rangle$ are Schwartz functions on $\mathsf{V}_{m,+}^{2}$)
@@
where $p\equiv(E(p),\vec{p})$, $k\equiv(E(k),\vec{k})$ with $E(p)=p^{0}=\sqrt{\vec{p}^{2}+m^{2}}$ and $E(k)=k^{0}=\sqrt{\vec{k}^{2}+m^{2}}$. The integrated functions of $\vec{p}$ and $\vec{k}$ are of Schwartz type in ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\equiv\mathsf{V}_{m,+}^{2}$ by construction. Thus, in particular, one can pass the $x$-derivatives under the symbol of integration. Let us discuss the proof of (d). First of all we observe that the integrals on both sides are finite since the integrated functions are Schwartz, in view of the bound (55), which is valid in two Minkowskian coordinate systems comoving with the reference frames $u_{\Sigma}$ and $u_{\Sigma^{\prime}}$ respectively. Define $J^{\nu}:=v_{\mu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle$. We observe that $J$ is smooth and conserved, $\partial_{\mu}J^{\mu}(x)=0$, by (c). Now consider the case where $\Sigma$ and $\Sigma^{\prime}$ are parallel. In this case we can use a common system of Minkowskian coordinates, where $\Sigma$ is the plane $x^{0}=0$ and $\Sigma^{\prime}$ is the analogous plane at $x^{0}=T>0$ (or vice versa). If $C_{r}\subset{\mathbb{R}}^{4}$ is the cylinder with axis $\vec{x}=0$ and basis $B_{r}:=\{\vec{x}\in{\mathbb{R}}^{3}\>|\>|\vec{x}|<r\}$ in $\Sigma$, we can use the divergence theorem for the compact cylinder $G_{r}\subset C_{r}$ bounded by $B_{r}$ and the analogous basis $B^{\prime}_{r}$ in $\Sigma^{\prime}$. We move on to compute the flux of $J$ across $\partial G_{r}$. This flux must be 0 since $J$ is conserved, taking the divergence theorem into account. The flux of $J$ on the lateral surface of $G_{r}$ is bounded by $cTr^{2}$, so that it tends to 0 as $r\to+\infty$ due to (55). The outward fluxes on the bases respectively tend to $\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma}d\Sigma$ and $\int_{\Sigma^{\prime}}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma^{\prime}}d\mu_{\Sigma^{\prime}}$ in view of the dominated convergence theorem. This proves the thesis for the case under consideration. It remains to consider the case where $\Sigma$ and $\Sigma^{\prime}$ are not parallel and thus meet in a 22-plane. Let us arrange a Minkowskian reference frame with the origin contained in this intersection and $\Sigma$ described by the plane $x^{0}=0$ in ${\mathbb{R}}^{4}$. With a further spatial rotation of the coordinates if necessary, we can describe the 33-plane $\Sigma^{\prime}$ as the plane $x^{0}=\gamma x^{1}$ in ${\mathbb{R}}^{4}$ where $\gamma\in(0,1)$, since $\Sigma^{\prime}$ is spacelike. We assume that we deal with vectors $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ such that their (finite smooth) components $\Psi^{n},\Psi^{\prime n}$ have compact support in the corresponding spaces ${\mathbb{R}}^{3n}\equiv\mathsf{V}_{m,+}^{n}$. At this juncture, using the same proof of (b) in (19) of Theorem [CDM26] for the integrals (118), we have that the second bound in (b) is satisfied:
@@
Proof of Lemma 6.3. It is sufficient to apply Eq.(3) in [Few12] referring to the explicit computations presented in Sec.2.5 therein, specialized to the Minkowski spacetime and using as reference state the Poincaré invariant state we indicated by $\Omega$. (The only necessary hypothesis is that the two-point function of the considered normalized state vector $\Psi$ is such that ${\mathbb{M}}\times{\mathbb{M}}\ni(x,y)\mapsto\langle\Psi|\hat{\phi}(x)\hat{\phi}(y)\Psi\rangle-\langle\Omega|\hat{\phi}(x)\hat{\phi}(y)\Omega\rangle\in C^{\infty}({\mathbb{M}}\times{\mathbb{M}})$. For $\Psi\in{\mathfrak{S}}_{0}$ this fact is guaranteed by Proposition $4.$) In this case the partial differential operator $Q$ is the operator $u^{\mu}u^{\prime\nu}D_{\mu\nu}$ (47) treated as in the proof of Proposition 4.10: $u^{\mu}u^{\prime\nu}D_{\mu\nu}=c^{2}D^{\prime}_{00}-s^{2}D^{\prime}_{11}$